OpenStudy (anonymous):

on three consecutive passes a football team gains 7 yards, loses 28 yards and gains 11 yards. what never represents the total net yardage.

OpenStudy (mattfeury):

7 - 28 + 11

OpenStudy (anonymous):

ive got that so far. -21+11 so would it be 10?

OpenStudy (mattfeury):

-10. -21 + 11 = 11 - 21 = -10

OpenStudy (anonymous):

so total net yardage they got is -10?

OpenStudy (mattfeury):

yep. does that make sense?

OpenStudy (anonymous):

well dont i need to rearrange them?

OpenStudy (mattfeury):

addition follows the commutative property in that order doesn't matter: http://en.wikipedia.org/wiki/Commutative_property any order you add them in will be the same result.

myininaya (myininaya):

negative shows u lost yardage so -10 would make sense not 10

myininaya (myininaya):

u lost more than you gain

OpenStudy (anonymous):

ok i see what your saying.

OpenStudy (anonymous):

i was thinking of putting them 11+7-28. but that would be way off. I see where you got -10. thanks soo much